OpenStudy (anonymous):

If I am currently making a 92.19% and have to take a test that has a maximum student score of 320, what is the lowest score I can make to stay above a 90% ?

OpenStudy (reemii):

What is the weight that will be given to the "92.19" and the next score? is it 50-50? I think something is missing in the statement.

OpenStudy (anonymous):

This is what my grade book looks like, I know I worded it wrong, but I'm not sure how to tell from what I have here. http://24.media.tumblr.com/380de86fc168c9e8f1c790c573b2aadc/tumblr_mlxj4duMBV1sp7n2ho1_500.jpg

OpenStudy (anonymous):

If you can't see it, its says: student score: NA/720 percentage: 0 points possible: 320 points earned: 0 I haven't taken the test yet.

OpenStudy (reemii):

Oh, the number 720 is what we need.

OpenStudy (reemii):

You got some scores and it have added up to \(n/720\) which is \(92.19/100\) (92.19 percent). This means you had more or less \(n=664\). Now you will have a new test and obtain a new score: \(b/320\). You need to find \(b\) so that \((n+b)/(720+320)=\frac{90}{100}\)

OpenStudy (reemii):

\(b = 1040\frac{90}{100} - n = 936-664=272\). Maybe with the approximation (\(n=\)663,...) it is better to say "at least 273".

OpenStudy (anonymous):

Oh okay, thank you so much!

OpenStudy (reemii):

yw